OpenStudy (pld51):

A circle is defined by the equation x2 + y2 – x – 2y – = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (sugarpie):

what the name of the paper

OpenStudy (eliesaab):

Complete the sqaures

OpenStudy (eliesaab):

\[x^2 + y^2 -x -2 y=0\\ x^2 -x +\frac 1 4-\frac 1 4+y^2 -2y +1 -1=0\]

OpenStudy (sugarpie):

K let me think a min

OpenStudy (eliesaab):

You can finish it now

OpenStudy (sugarpie):

is there opeshions

OpenStudy (eliesaab):

It seems that he left

OpenStudy (sugarpie):

k

OpenStudy (amorfide):

you must complete the square so for example \[x^{2}+2x=(x+\frac{ 2 }{ 2 })^{2}- (\frac{ 2 }{ 2 })^{2}\] completing the square for an expression ax^2 + bx you have \[ax^{2}+bx=a(x^{2}+\frac{ b }{ a }x)=a((x+\frac{ b }{ 2a })^{2}-(\frac{ b }{ 2a })^{2})\] so where you have x2 + y2 – x – 2y – = 0. pair up the expressions of x and y x2 -x +y2 -2y=0 then we complete the square for the x terms x2-x a=1 b=-1 so substitute that into what i gave you then you complete the square for the y terms y2-2y a=1 b=-2 and then you make it all equal to 0 then you will move your constant to the right side \[(x+a)^{2}+(y+b)^{2}=r^{2}\] with a centre (-a,-b), radius r

OpenStudy (eliesaab):

\[ x^2 + y^2 -x -2 y=0\\ x^2 -x +\frac 1 4-\frac 1 4+y^2 -2y +1 -1=0\\ (x-1/2)^2 + (y-1)^2= 1/4 +1 \]

OpenStudy (eliesaab):

You are done now

OpenStudy (sugarpie):

the answer is 2 units If x² + ax + y² + by + c = 0 then (x + ½a)² + (y + ½b)² = ¼(a² + b² - 4c) So, for the general form, the center is (-½a, -½b) and the radius is ½√(a² + b² - 4c) Here a=-1, b=-2, and c=-11/4, so the center is (½, 1) and the radius is ½√(1 + 4 + 11) which is 2 units, hence the

OpenStudy (sugarpie):

did that help